# First is Not Really Better Than Last: Evaluating Layer Choice and Aggregation Strategies in Language Model Data Influence Estimation

### v1

@@ Appendix A Experimental setup details
Checkpoints preparation adds LoRA modules [Dettmers-2023] to query and value projections in every attention layer, compresses embeddings to the tokens participating in the GLUE dataset splits. The embedding compression greatly speeds up influence computations based on gradients.
@@
Stage 2. Models are fine-tuned on noisy datasets for 10 epochs from the initial checkpoint. Validation loss and accuracy are tracked to select the checkpoint for influence estimation. Following [Garima-2020] and preliminary experiments, we use the checkpoint with the lowest validation loss in the following stages.
@@ Appendix F Best test set accuracy after filtering
The following tables 4,5,6 present the ranking of filtering methods based on different influence scoring and aggregation for 3 models under consideration. Rows are ordered according to mean rank from the best method to the worst. Two baselines are Random (dropping 30% uniformly) and Full (removing all noise (20%) and 10% of random clean samples). A good filtering method should outperform the Random, the lower bound, and approach the Full, upper bound. The tables highlight the best non-baseline methods for each GLUE dataset. AUC measure is specified for the NDR curve; a high value indicates skew of noise towards the beginning of the influence range.
@@ Appendix G Proof of Cancellation Effect Theorem
i.e. the separation between noisy and clean samples is strictly larger under $I_{\theta,\omega}$, thereby showing that contrary to the claim of [Yen-2022], the inclusion of weights with high cancellation can improve influence estimates.
@@
Remark. This shows that contrary to the claim of [Yen-2022], inclusion of weights with high cancellation can improve discrimination in influence scoring, although the improvement depends on the choice of validation sample. Indeed, there are also cases where $\Delta I_{\theta,\omega}<\Delta I_{\theta}$, so the proposition in [Yen-2022] cannot be concluded in full generality.
@@ Appendix I Noise influence distribution on model layers
An influence function is not useful for anomaly detection if it is closer to a uniform distribution. For selected methods, most mislabeled samples are concentrated at the beginning. However, some layers and methods have a high influential noise level. For instance, TracIn demonstrates relatively high spikes on both sides across configurations. TracIn on WE and CL has a high percentage of influential noise. Generally, a performant attribution should strive to minimize the noise entropy, producing unimodal but not uniform distributions. The figures 192021 present how mislabeled samples are spread in the influence scores. In the experiments, we discard the first two quantiles.

### v2

@@ 6 Conclusion
7 Reproducibility Statement
We have open-sourced our code and implementation for our experiments: https://github.com/dvitel/nn-infl. To support reproducibility, multiple runs were executed with fixed random seeds specified in the accompanying shell scripts, and determinism was enabled in the underlying pytorch framework. All experiments were conducted on an Ubuntu server with 8x NVIDIA A100s (40GB VRAM/GPU).
@@ Appendix A Experimental setup details
Checkpoints preparation adds LoRA modules [Dettmers et al., 2023] to query and value projections in every attention layer, compresses embeddings to the tokens participating in the GLUE dataset splits. The embedding compression greatly speeds up influence computations based on gradients.
@@
Stage 2. Models are fine-tuned on noisy datasets for 10 epochs from the initial checkpoint. Validation loss and accuracy are tracked to select the checkpoint for influence estimation. Following [Pruthi et al., 2020] and preliminary experiments, we use the checkpoint with the lowest validation loss in the following stages.
@@ Appendix F Best test set accuracy after filtering
The following tables 4,5,6 present the ranking of filtering methods based on different influence scoring and aggregation for 3 models under consideration.
@@
Rows are ordered according to mean rank from the best method to the worst. Two baselines are Random (dropping 30% uniformly) and Full (removing all noise (20%) and 10% of random clean samples). A good filtering method should outperform the Random, the lower bound, and approach the Full, upper bound. The tables highlight the best non-baseline methods for each GLUE dataset. AUC measure is specified for the NDR curve; a high value indicates skew of noise towards the beginning of the influence range.
@@ Appendix G Proof of Cancellation Effect Theorem
i.e. the separation between noisy and clean samples is strictly larger under $I_{\theta,\omega}$, thereby showing that contrary to the claim of [Yeh et al., 2022], the inclusion of weights with high cancellation can improve influence estimates.
@@
Remark. This shows that contrary to the claim of [Yeh et al., 2022], inclusion of weights with high cancellation can improve discrimination in influence scoring, although the improvement depends on the choice of validation sample. Indeed, there are also cases where $\Delta I_{\theta,\omega}<\Delta I_{\theta}$, so the proposition in [Yeh et al., 2022] cannot be concluded in full generality.
@@ Appendix I Noise influence distribution on model layers
An influence function is not useful for anomaly detection if it is closer to a uniform distribution. For selected methods, most mislabeled samples are concentrated at the beginning. However, some layers and methods have a high influential noise level. For instance, TracIn demonstrates relatively high spikes on both sides across configurations. TracIn on WE and CL has a high percentage of influential noise. Generally, a performant attribution should strive to minimize the noise entropy, producing unimodal but not uniform distributions.
The following figures 202122 present how mislabeled samples are spread in the influence scores. In the experiments, we discard the first two quantiles.
@@
Appendix J Influential data identification on autoregressive datasets
Datasets. To further validate the RQ2 (first is not better than last), we conduct the experiment on the autoregressive datasets introduced in Kwon et al. [2024]: Grammar, Math, and Math (With Reasoning). Unlike the settings used in the main text, where detrimental samples were synthetically injected, these datasets contain semantically related instruction categories. In this context, a desirable behavior is that training instructions exert a stronger influence on test samples from the same semantic category. In this way, we assess how well influence scores recover meaningful relational structure rather than merely detecting artificial noise. Such scenarios are highly relevant for practical applications, including knowledge probing, debugging data pipelines, and detecting harmful or misaligned training examples.
New methods. Additionally, we extend our experiments to incorporate more recent gradient- and activation-based influence estimation techniques. We report the following results for Outlier Gradient Chhabra et al. [2025], Kronfluence (EKFACGrosse et al. [2023], and RepSim Li et al. [2025]. We detect gradient outliers by scoring training samples with OneClassSVM using an RBF kernel on a per-semantic-category basis. EKFAC scoring combines forward and backward signals to estimate influence. RepSim is based purely on hidden representations measured after each attention layer either on last sample token or averaged across all tokens of the sample (mean).
Models. The results are provided for Qwen2.5-1.5B (lr=3e-4) and Mistral-7B-v0.3 (lr=1e-3).
Metrics. We use the evaluation metrics, AUC and Recall, from the original work Kwon et al. [2024]. Higher values indicate bettser alignment between the sample ordering by the influence technique and the actual semantic category of the test samples.
Conclusions. Figures 2324 present the AUC and Recall across layers for Qwen, and  2526 - for Mistral. We observe that RQ2 conclusion still holds for new datasets.
First, early layers consistently exhibit weaker performance in AUC and Recall for both models across methods. Similarly, the last attention layers show a drop in AUC-Recall in the majority of cases. Unfortunately, the best-performing layers vary with model-dataset-method, but generally, we observe the performance maximum for 20-24 layers.
Second, for both Qwen and Mistral, the LoRA B module (attached to value projections) delivers the strongest performance across most configurations. Our practical recommendation from consistent observations across all evaluated models and datasets is to rely on influence scores computed on these modules. These weights yield the highest NDR, AUC, Recall, and downstream task performance.
Figure 23: Qwen 2.5 1.5B AUC on influence datasets.
Figure 24: Qwen 2.5 1.5B Recall on influence datasets.
Figure 25: Mitral 7B AUC on influence datasets.
Figure 26: Mistral 7B Recall on influence datasets.
Appendix K Sample influence variation across model layers
We provide the following visualization of the sample influence ranking trends as additional interpretation of why middle layers have better performance. Figures 27 and  28 present these ranking, where highest rank means highest influence score, for Roberta-Large and Qwen-2.5 1.5B, respectively, and three “kinds” of samples:
  • Top Ranked Noise Sample - the noisy sample with the highest influence across layers;
  • Avg Noise - the averaged rank of all noisy samples per layer;
  • Avg Benign - the averaged rank of “clean” samples per layer.
Figure 27: Roberta-Large sample influence rank variation with layer. Mean and 95% confidence intervals are presented for 10 random seeds.
Conclusion. We observe that the average noise influence rank drops with layer, i.e. noise becomes less influential. At the same time, the benign samples maintain the influence rank almost at same level. For Roberta-Large, the biggest difference between average ranks is observed at the last influence layers. This corresponds to the downstream performance maximum (Figure 6), where the best accuracy is also observed on later attention layers. Similarly, for Qwen-2.5 1.5B, the noise average rank has minimum on layers 15-20, where the model has also the best downstream performance (Table 3).
At the same time, for Qwen, in many cases, the top ranked noisy sample maintain high values of influence when the average noise rank has minimum, signifying the presence of the noise outliers that may affect the performance.
Figure 28: Qwen-2.5 1.5B sample influence rank variation with layer. Mean and 95% confidence intervals are presented for 10 random seeds.
Appendix L The effect of varying the number of votes in positional voting
Formula 10 contains the hyperparameter $k$, the number of votes that each test sample assigns to training samples. The reasonable question is how the noise detection rate changes with $k$ and what $k$ would correspond to the best performance? The following figure 29 presents the NDR for $k$ in range $[10,100]$ with step $10$, for different layers and per NN module of Qwen-2.5 1.5B. The NDR value is averaged across datasets and random seeds.
Figure 29: Qwen-2.5 1.5B NDR variation for different Vote $k$ and layer number.
Conclusion. NDR values show a clear dependence on the choice of $k$. The range $k\in[10,50]$ yields the strongest performance for most settings. For DataInf and TracIn, NDR decreases monotonically beyond this range, suggesting reduced discriminative power at larger subset sizes. Cosine shows the opposite pattern, reaching its minimum at $k\approx 50$ and improving for larger $k$. We use $k=30$ (filtering threshold) in all main experiments. Determining how the optimal $k$ scales with the level of injected noise remains an open question.
Appendix M Llama-3.2 1B influence score performance discussion
In Section 5.2, we compare the training data filtering by influence attribution with random baseline. We found, that particularly for LLaMA-3.2 1B, the influence functions fail to outperform random filtering. In this section, we verify if the best by NDR module (Figure 15), LoRA B attached to value projection, outperforms the baseline.
The artifacts observed for LLaMA-3.2-1B likely arise from its reduced depth (16 attention layers) relative to larger models (Roberta-Large - 24, Qwen-2.5 1B - 28, Mistral 7B - 32). The variation in best-performing layers across architectures indicates that reliable influence attribution requires not only well-chosen probes but also effective layer localization. This layer dependence parallels findings in other domains where the success of model interventions hinges on identifying the most relevant layers.
The following table 7 presents the accuracy on GLUE datasets after influence filtering by Value B score of 8th and 9th attention layers (best according to NDR). Results include Mean, Rank, and Vote aggregations.
Table 7: Llama-3.2 1B downstream performance of 8th and 9th attention layer influence scores measured for LoRA modules B that are attached to value projections.
Method Agg Layer Rank QNLI MRPC SST2 QQP COLA MNLI RTE STSB
Full 1.20 $\pm$ 0.80 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.6 $\pm$ 1.9 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 2.1 $\pm$ 1.0 1.0 $\pm$ 0.0
Cosine Vote v-b-9 6.80 $\pm$ 4.10 6.0 $\pm$ 4.5 8.9 $\pm$ 4.1 12.8 $\pm$ 2.2 5.1 $\pm$ 2.3 4.6 $\pm$ 1.5 7.3 $\pm$ 3.5 6.6 $\pm$ 3.7 3.5 $\pm$ 1.3
Cosine Vote v-b-8 6.90 $\pm$ 4.40 5.2 $\pm$ 2.0 11.2 $\pm$ 5.1 12.0 $\pm$ 3.4 5.5 $\pm$ 2.2 6.4 $\pm$ 2.9 5.2 $\pm$ 3.3 7.2 $\pm$ 5.0 2.8 $\pm$ 1.0
DataInf Vote v-b-8 9.20 $\pm$ 4.50 13.1 $\pm$ 4.9 7.1 $\pm$ 5.1 9.8 $\pm$ 2.5 8.0 $\pm$ 4.3 6.4 $\pm$ 3.7 11.6 $\pm$ 3.4 10.6 $\pm$ 4.8 7.0 $\pm$ 2.6
Random 9.40 $\pm$ 6.00 8.3 $\pm$ 4.8 10.6 $\pm$ 7.1 13.3 $\pm$ 5.1 3.8 $\pm$ 2.1 6.8 $\pm$ 3.2 14.5 $\pm$ 2.2 14.4 $\pm$ 6.5 3.6 $\pm$ 1.4
DataInf Vote v-b-9 9.80 $\pm$ 5.20 16.0 $\pm$ 4.0 5.5 $\pm$ 5.0 8.8 $\pm$ 3.3 7.6 $\pm$ 4.8 5.2 $\pm$ 3.0 15.0 $\pm$ 3.3 11.0 $\pm$ 3.1 9.2 $\pm$ 3.3
TracIn Vote v-b-8 9.90 $\pm$ 4.80 14.6 $\pm$ 3.0 7.6 $\pm$ 6.6 8.6 $\pm$ 4.4 9.4 $\pm$ 4.0 7.2 $\pm$ 2.5 11.8 $\pm$ 4.3 11.2 $\pm$ 5.5 8.9 $\pm$ 3.3
Cosine Rank v-b-8 10.40 $\pm$ 5.10 6.6 $\pm$ 3.7 15.8 $\pm$ 3.3 8.1 $\pm$ 5.4 12.2 $\pm$ 3.5 14.4 $\pm$ 2.5 4.6 $\pm$ 2.8 9.2 $\pm$ 5.3 12.2 $\pm$ 1.7
Cosine Rank v-b-9 10.60 $\pm$ 5.00 8.8 $\pm$ 4.4 12.7 $\pm$ 5.4 6.8 $\pm$ 4.1 12.0 $\pm$ 3.4 12.4 $\pm$ 3.2 7.6 $\pm$ 4.1 8.4 $\pm$ 4.7 16.4 $\pm$ 3.9
TracIn Vote v-b-9 10.60 $\pm$ 5.60 17.4 $\pm$ 3.2 5.2 $\pm$ 4.4 9.4 $\pm$ 4.2 9.2 $\pm$ 4.8 7.0 $\pm$ 3.0 15.2 $\pm$ 4.5 12.0 $\pm$ 4.9 9.0 $\pm$ 4.0
Cosine Mean v-b-8 10.70 $\pm$ 5.20 10.5 $\pm$ 5.6 15.0 $\pm$ 4.6 5.3 $\pm$ 3.3 13.4 $\pm$ 3.3 15.2 $\pm$ 3.0 9.5 $\pm$ 4.2 7.4 $\pm$ 5.6 9.2 $\pm$ 3.1
DataInf Mean v-b-8 11.10 $\pm$ 5.60 10.4 $\pm$ 4.3 12.6 $\pm$ 5.3 13.6 $\pm$ 6.1 9.9 $\pm$ 6.6 8.8 $\pm$ 5.2 12.6 $\pm$ 5.1 9.3 $\pm$ 6.7 11.8 $\pm$ 5.5
Cosine Mean v-b-9 11.40 $\pm$ 5.40 10.2 $\pm$ 2.4 13.4 $\pm$ 4.2 5.0 $\pm$ 3.7 13.6 $\pm$ 3.2 16.5 $\pm$ 2.9 12.5 $\pm$ 5.5 6.2 $\pm$ 5.4 13.6 $\pm$ 4.7
DataInf Rank v-b-8 11.70 $\pm$ 4.60 8.6 $\pm$ 4.4 11.4 $\pm$ 4.5 10.6 $\pm$ 4.3 16.4 $\pm$ 1.6 16.2 $\pm$ 2.5 7.4 $\pm$ 3.0 13.0 $\pm$ 3.8 10.1 $\pm$ 3.4
TracIn Rank v-b-8 12.70 $\pm$ 4.80 10.7 $\pm$ 6.0 12.2 $\pm$ 3.1 12.0 $\pm$ 3.8 17.0 $\pm$ 2.3 16.4 $\pm$ 3.1 7.0 $\pm$ 4.0 14.0 $\pm$ 4.3 12.6 $\pm$ 3.1
TracIn Mean v-b-8 12.80 $\pm$ 5.70 13.2 $\pm$ 4.8 13.6 $\pm$ 5.5 16.6 $\pm$ 3.4 12.0 $\pm$ 7.0 8.0 $\pm$ 5.2 14.3 $\pm$ 5.5 10.4 $\pm$ 6.0 14.6 $\pm$ 5.1
DataInf Rank v-b-9 12.80 $\pm$ 4.50 10.8 $\pm$ 4.9 11.7 $\pm$ 4.1 12.4 $\pm$ 4.9 15.1 $\pm$ 1.7 15.8 $\pm$ 3.0 8.4 $\pm$ 4.5 12.6 $\pm$ 5.0 15.2 $\pm$ 3.0
DataInf Mean v-b-9 13.10 $\pm$ 5.60 11.6 $\pm$ 6.1 12.4 $\pm$ 3.9 16.0 $\pm$ 7.3 9.0 $\pm$ 4.9 10.6 $\pm$ 4.9 16.2 $\pm$ 5.2 13.2 $\pm$ 4.7 16.0 $\pm$ 3.6
TracIn Rank v-b-9 14.00 $\pm$ 4.90 11.5 $\pm$ 5.8 11.1 $\pm$ 3.3 12.2 $\pm$ 5.9 17.2 $\pm$ 3.3 17.9 $\pm$ 2.0 10.5 $\pm$ 4.3 15.2 $\pm$ 4.7 16.5 $\pm$ 2.7
TracIn Mean v-b-9 14.80 $\pm$ 5.30 15.6 $\pm$ 5.7 11.2 $\pm$ 5.1 15.2 $\pm$ 7.1 12.6 $\pm$ 5.9 13.0 $\pm$ 4.3 17.8 $\pm$ 4.2 16.0 $\pm$ 3.1 17.0 $\pm$ 3.4
We observe that Cosine scores (gradient similarities) outperform Random with Vote aggregation ($k=20$) on selected modules. The differences are statistically significant. Wilcoxon test p-value is 0.0072 for configuration (Cosine, Vote, v-b-9) and Random; p-value is 0.0085 for (Cosine, Vote, v-b-8) and Random. At the same time, (DataInf, Vote, v-b-8) and Random have the p-value 0.47, signifying that the null hypothesis of their similarity cannot be rejected. At the same time, Mean aggregation cannot outperform the baseline even on the best NDR layers.
Conclusion. With the new Vote aggregation strategy, influence scoring outperforms the baselines on locations that perform best according to NDR. Our practical recommendation for LLaMA-3.2-1B, and consistently across the evaluated models, is to rely on influence scores computed on the LoRA B value projection modules. These weights yield strong NDR performance also on autoregressive datasets (Appendix J).